How many formula units does 200.0 g of calcium carbonate represent
1
Expert's answer
2020-04-18T07:07:40-0400

200.0 g calcium carbonate contain 200.0/100.09 = 1.998 mol and number of formula units is 1.998*6.02*1023 = 1.203*1024.